1. There are two potential consumers for a public good. Consumer 1 has inverse demand P 1 = 30 − Q, Consumer 2 has inverse demand P 2 = 60 − 3 Q, where Q is the level of provision of the public good. The marginal cost of providing the public good is 30. What is the optimal level of provision of the public good?

30

0

25

15

2. Production of widgets creates pollution: a negative externality. The marginal private cost of producing widgets is constant at $10. The marginal cost of pollution is increasing, M E C ( Q ) = Q. Inverse demand in the market for widgets is given by P = 110 − Q. What is the optimal Pigouvian tax in this market?

Group of answer choices

10

50

110

60

3. There are three potential consumers for a public good. Each consumer has an inverse demand for the public good given by P = 200 − 2 Q, where Q is the level of provision of the public good. The marginal cost of providing the public good is 300. What is the optimal level of provision of the public good?

Group of answer choices

25

100

0

50
Answer #1

1. 15
(Optimal level is determined where P1 + P2 = MC
So, 30 - Q + 60 - 3Q = 90 - 4Q = 30
So, 4Q = 90 - 30 = 60
So, Q = 60/4 = 15)

2. 50
(P = MC + MEC
So, 110 - Q = 10 - Q
So, Q = Q = 2Q = 110 - 10 = 100
So, Q = 100/2 = 50
Pigouvian tax = MEC = Q = 50)

3. 100
(Optimal level is determined where 3P = MC
So, 3(200-Q) = 600 - 3Q = 300
So, 3Q = 600-300 = 300
So, Q = 300/3 = 100)